OpenStudy (anonymous):

x+3/3=y+2/2 then x/3=?

OpenStudy (anonymous):

this is confusing to me. my textbook doesnt explain it too well.

OpenStudy (anonymous):

you've been here for a month now. try using "TEX" or the equation button to type your equation in. Your question is, (I can safely say) confusing to evrybody

OpenStudy (anonymous):

i dont have TEX or equation button. and excuse me for being "here for a month" and not understanding.

OpenStudy (anonymous):

and not understanding all the "right ways" to enter

OpenStudy (anonymous):

just below the reply box, you'd find a blue button that says equation. please enter the equation so we know what it exactly is.

OpenStudy (anonymous):

\[\frac{ x+3 }{ 3 }=\frac{ y+2 }{ 2 } then \frac{ x }{ 3 }=?\]

OpenStudy (anonymous):

ok \[{x\over3}+{3\over3}={y\over2}+{2\over2}\\ \implies{x\over3}+\cancel{1}={y\over2}+\cancel{1}\\ \implies{x\over3}={y\over2}\]

OpenStudy (anonymous):

does the question have any further information like a point, or something?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then that is your answer!

OpenStudy (anonymous):

I was making it harder than it is. I thought that was too obvious. Thanks for explaining by taking the 1s out. That makes sense.

OpenStudy (anonymous):

surething. see, It will be very useful if the asker and asked help eachother.